# CT-PatchTST: Channel-Time Patch Time-Series Transformer for Long-Term Renewable Energy Forecasting

### v1

@@ Abstract
Accurately predicting renewable energy output is crucial for the efficient integration of solar and wind power into modern energy systems. This study develops and evaluates an advanced deep learning model, Channel-Time Patch Time-Series Transformer (CT-PatchTST), to forecast the power output of photovoltaic and wind energy systems using annual offshore wind power, onshore wind power, and solar power generation data from Denmark. While the original Patch Time-Series Transformer(PatchTST) model employs a channel-independent (CI) approach, it tends to overlook inter-channel relationships during training, potentially leading to a loss of critical information. To address this limitation and further leverage the benefits of increased data granularity brought by CI, we propose CT-PatchTST. This enhanced model improves the processing of inter-channel information while maintaining the advantages of the channel-independent approach. The predictive performance of CT-PatchTST is rigorously analyzed, demonstrating its ability to provide precise and reliable energy forecasts. This work contributes to improving the predictability of renewable energy systems, supporting their broader adoption and integration into energy grids.
@@
power generation prediction, renewable energy, self-attention, long time series, time series forecasting
@@
I Introduction
Clean energy has become a cornerstone of global energy systems, playing a pivotal role in advancing sustainability and combating climate change. With growing concerns about the environmental impact of fossil fuels, renewable energy sources, particularly wind and solar power, have seen widespread adoption worldwide.[1] For example, The EU added 11,791.4 MW of installed wind capacity in 2014, accounting for 9.16% of the total installed capacity.[2] These two forms of renewable energy not only account for a significant share of global clean energy production but also serve as key drivers in shaping sustainable energy policies and practices. As renewable energy technologies continue to evolve, accurately forecasting wind and solar power output has emerged as a critical task for optimizing energy management and ensuring their seamless integration into modern energy grids.
Forecasting power generation is essential for maintaining a stable balance between energy supply and demand, enhancing the efficiency of energy storage systems, and mitigating risks of grid instability. Accurate predictions enable energy providers, grid operators, and policymakers to plan energy distribution, storage, and operational strategies more effectively. However, renewable energy forecasting poses unique challenges due to its dependence on external and often unpredictable natural factors.[3] For instance, wind power generation is influenced by variables such as wind speed, direction, and atmospheric temperature, while solar power generation depends on sunlight intensity, cloud cover, and seasonal variations. These factors introduce nonlinear and dynamic patterns in power generation, further compounded by diurnal and seasonal cycles, making reliable forecasting a complex yet vital task.
Traditional methods for renewable energy forecasting, such as statistical models and physical simulations, have provided a strong foundation for the field but face significant limitations in handling the intricate temporal dependencies and nonlinear relationships inherent in renewable energy data. Statistical models like ARIMA, while effective for short-term forecasting, often struggle to capture long-term dependencies and adapt to abrupt changes in data patterns.[4][5] On the other hand, physical simulations require extensive domain-specific knowledge and computational resources, limiting their scalability in dynamic real-world scenarios. These challenges underscore the urgent need for advanced and flexible approaches to address the complexities of renewable energy forecasting.
Recent advancements in deep learning have provided powerful tools for overcoming these challenges, like filtering-based methods[6][7] and support vector machine[8]. For example, Justin used the combination of decision trees and support vector regression to decrease the prediction time. [9] Among these methods, the transformer[10] and Patch Time-Series Transformer (PatchTST)[11] has demonstrated remarkable potential in capturing temporal dependencies within multivariate time-series data. However, the original PatchTST employs a channel-independent approach, which, while effective for managing data granularity, often neglects critical inter-channel relationships. This limitation can result in the loss of valuable contextual information essential for accurate forecasting.[12] [13]To address these shortcomings, we propose an enhanced model, the Channel-Time Patch Time-Series Transformer (CT-PatchTST). Fig. 1 provides an overview of the model architecture. CT-PatchTST builds upon the strengths of PatchTST by introducing channel attention mechanisms that effectively capture and integrate inter-channel dependencies while retaining the benefits of channel-independent modeling. The key contributions of this paper are as follows:
@@ II Related Work
Time-series forecasting has long been a critical area of study, evolving from traditional statistical models to sophisticated deep learning techniques. Early approaches, such as the AutoRegressive Integrated Moving Average (ARIMA) model, laid the groundwork for this field. [14] ARIMA’s design effectively handles stationary and non-stationary data by incorporating autoregressive, moving average, and differencing components. While powerful for short-term predictions, ARIMA falls short in addressing long-term dependencies and complex non-linear relationships.
@@
Exponential smoothing methods, like the Holt-Winters model, have also been extensively used for capturing trends and seasonality. By assigning diminishing weights to older observations, these models perform well for data with seasonal variations. However, their inability to model multivariate and non-linear patterns limits their application. Gaussian Process Regression (GPR), a probabilistic framework, offers another perspective by modeling time series as distributions over functions, providing both predictions and uncertainty estimates. However, GPR struggles to scale efficiently with large datasets.
@@
To overcome the limitations of traditional methods, deep learning models have emerged as powerful tools for time-series forecasting. Among these, Long Short-Term Memory (LSTM) networks are particularly notable for their ability to model long-term dependencies. By utilizing a cell state to retain relevant information and gate mechanisms to regulate data flow, LSTMs effectively address the vanishing gradient problem.[15] For example, in applications like oil production forecasting, LSTM-based frameworks have been shown to outperform traditional methods significantly, demonstrating their adaptability to complex and non-linear temporal dynamics.
@@
Building on the advancements made by LSTMs, Seq2Seq models have brought further innovation to time-series forecasting. Initially developed for natural language processing tasks, Seq2Seq models employ an encoder-decoder structure where the encoder compresses the input sequence into a context vector, and the decoder generates predictions based on this representation. This architecture enables Seq2Seq models to handle variable-length input and output sequences, making them suitable for time-series prediction tasks with diverse temporal dynamics.[16] However, their reliance on fixed-length context vectors can result in information bottlenecks, particularly for long sequences. The introduction of attention mechanisms has helped mitigate this issue by enabling the model to focus on the most relevant parts of the input sequence dynamically. Despite their computational intensity and scaling challenges, Seq2Seq models have demonstrated notable success in capturing complex temporal dependencies, paving the way for more advanced architectures.
@@
Recently, the introduction of Transformer-based architectures has revolutionized time-series forecasting, with the Patch Time-Series Transformer (PatchTST) emerging as a notable example. PatchTST segments input sequences into non-overlapping patches, reducing sequence length and enabling efficient parallel processing while maintaining temporal granularity. Its self-attention mechanism dynamically assesses the importance of various time steps, making it highly effective in capturing both local and global temporal patterns. This patch-based approach not only enhances scalability but also improves model generalization across diverse datasets. Compared to Seq2Seq models, although it requires careful hyperparameter tuning and considerable computational resources during training, PatchTST has demonstrated superior performance in long-horizon forecasting tasks.
@@ III-B Reversible Instance Normalization
Reversible Instance Normalization (RevIN) is designed to mitigate the distribution shift between training and testing data in time-series applications. We apply this technique to normalize the raw data before turning it into patches. This technique operates on each time-series instance $x^{(i)}$, where $x^{(i)}$ represents the sequence of values for the $i$-th feature or channel in the dataset, with $i=1,2,\dots,M$. Each $x^{(i)}$ consists of a series of temporal values $[x^{(i)}_{1},x^{(i)}_{2},\dots,x^{(i)}_{L}]$.
@@ III-D Channel-Time Transformer Encoder
where $\mathbf{Z}_{\text{channel}}^{m}\in\mathbb{R}^{P\times N}$ represents the features for channel $m$, and $\mathbf{W}_{h}^{Q,\text{time}},\mathbf{W}_{h}^{K,\text{time}},\mathbf{W}_{h}^{V,\text{time}}\in\mathbb{R}^{P\times d_{k}}$ are trainable parameters. The time attention output is computed as:
@@ IV-A Dataset
We evaluate our model using a dataset that comprises measured time series data of renewable energy production, electricity consumption, and energy exchange with neighboring countries and continents, all recorded at an hourly resolution.[17] The dataset spans the period from 2014 to 2019 and contains a total of 499,993 entries. Specifically, we focus on predicting three key variables: ’OffshoreWindPower,’ ’OnshoreWindPower,’ and ’SolarPowerProd.’ These variables provide valuable insights into renewable energy production patterns, making the dataset a robust benchmark for assessing the performance of our model in multivariate time series forecasting tasks.
@@
TABLE II: Full results on the multivariate forecasting task. We used a look-back window of length 336 for all datasets, and forecasting windows $h\in\{96,192,336,720\}$. The best results are highlighted in bold, and the second-best results are underlined.
Forecasting Horizon (h) CT-PatchTST-512 CT-PatchTST PatchTST Seq2Seq
Metric MSE MAE MSE MAE MSE MAE MSE MAE
96 0.0060 0.0515 0.0061 0.0504 0.0066 0.0540 0.0187 0.0561
192 0.0060 0.0505 0.0063 0.0502 0.0070 0.0548 0.0208 0.1190
336 0.0061 0.0509 0.0071 0.0532 0.0087 0.0595 0.1573 0.2270
720 0.0077 0.0579 0.0093 0.0644 0.0114 0.0683 0.1940 0.3510
@@
Results As shown in Table II, our model shows the best performance. This demonstrates that applying attention to channel data prior to the processing of multivariate sequences enhances the inter-channel relationships, thereby improving prediction accuracy.
@@ IV-D Ablation Study
In this section, we conducted two comprehensive ablation experiments to thoroughly examine the influence of varying patch lengths and prediction horizons on the performance of CT-PatchTST and to compare its effectiveness against the baseline PatchTST model. The first experiment focused on evaluating the impact of altering patch lengths, which were adjusted within a range from 4 to 40, to understand how the model’s ability to capture local and global patterns in multivariate time series data changes with different granularities. The second experiment explored the effect of varying prediction horizons, ranging from 96 to 720, to assess how well the models adapt to forecasting tasks of different temporal scales.

### v3

@@ Abstract
Accurate forecasting of renewable energy generation is fundamental to enhancing the dynamic performance of modern power grids, especially under high renewable penetration. This paper presents Channel-Time Patch Time-Series Transformer (CT-PatchTST), a novel deep learning model designed to provide long-term, high-fidelity forecasts of wind and solar power. Unlike conventional time-series models, CT-PatchTST captures both temporal dependencies and inter-channel correlations—features that are critical for effective energy storage planning, control, and dispatch. Reliable forecasting enables proactive deployment of energy storage systems (ESSs), helping to mitigate uncertainties in renewable output, reduce system response time, and optimize storage operation based on location-specific flow and voltage conditions. Evaluated on real-world datasets from Denmark’s offshore wind, onshore wind, and solar generation, CT-PatchTST outperforms existing methods in both accuracy and robustness. By enabling predictive, data-driven coordination of ESSs across integrated source–grid–load–storage systems, this work contributes to the design of more stable, responsive, and cost-efficient power networks.
@@
Energy forecasting, energy storage systems, dispatch optimization, transformer models, grid dynamic enhancement, smart grid control, multivariate time series.
[…]
=-21pt
I Introduction
Renewable energy has become a cornerstone of modern power systems, driven by the urgent need for sustainable and low-carbon energy solutions. However, its inherent variability poses serious challenges to real-time power balancing and grid stability. Energy storage systems (ESSs) have emerged as key enablers to mitigate such volatility, by shifting renewable output temporally and enabling dynamic response to grid conditions. Accurate forecasting of wind and solar generation is thus essential not only for day-ahead scheduling but also for the optimal planning, control, and dispatch of ESSs within integrated source–grid–load–storage frameworks. This paper introduces CT-PatchTST, a novel transformer-based forecasting model designed to meet these practical demands by enhancing the temporal and structural understanding of multivariate renewable datasets.
Forecasting power generation is essential for maintaining a stable balance between energy supply and demand, enhancing the efficiency of energy storage systems, and mitigating risks of grid instability. Accurate predictions enable energy providers, grid operators, and policymakers to plan energy distribution, storage, and operational strategies more effectively. However, renewable energy forecasting poses unique challenges due to its dependence on external and often unpredictable natural factors [1]. For instance, wind power generation is influenced by variables such as wind speed, direction, and atmospheric temperature, while solar power generation depends on sunlight intensity, cloud cover, and seasonal variations. These factors introduce nonlinear and dynamic patterns in power generation, further compounded by diurnal and seasonal cycles, making reliable forecasting a complex yet vital task.
Traditional methods for renewable energy forecasting, such as statistical models and physical simulations, have provided a strong foundation for the field but face significant limitations in handling the intricate temporal dependencies and nonlinear relationships inherent in renewable energy data. Statistical models like ARIMA, while effective for short-term forecasting, often struggle to capture long-term dependencies and adapt to abrupt changes in data patterns [2, 3]. On the other hand, physical simulations require extensive domain-specific knowledge and computational resources, limiting their scalability in dynamic real-world scenarios. These challenges underscore the urgent need for advanced and flexible approaches to address the complexities of renewable energy forecasting.
Recent advances in machine learning, particularly deep learning, have provided a variety of powerful tools to address complex and data-intensive challenges. Traditional approaches such as filtering-based methods and support vector machines remain valuable in many scenarios [4, 5, 6]. In parallel, deep learning has been increasingly applied in areas such as infrastructure analysis, image-based prediction, and sequential decision-making [7, 8, 9]. Representative architectures, such as Convolutional Neural Networks [10], Recurrent Neural Networks [11], and Transformers [12], have significantly improved the capacity to model complex spatial and temporal relationships [13, 14, 15, 16, 17]. The emergence of large language models has further expanded the capabilities of learning systems, enabling new possibilities in knowledge extraction, interpretation, and task automation [18, 19, 20, 21, 22, 23]. These developments are underpinned by ongoing theoretical progress in statistical estimation and optimization, which continue to enhance the generalization and scalability of learning-based models [24, 25]. Among these methods, the Patch Time-Series Transformer (PatchTST) [26] has demonstrated remarkable effectiveness in modeling temporal dependencies in multivariate time-series data. However, the original PatchTST employs a channel-independent approach, which, while effective for managing data granularity, often neglects critical inter-channel relationships. This limitation can result in the loss of valuable contextual information essential for accurate forecasting [27, 28].To address these shortcomings, we propose an enhanced model, the Channel-Time Patch Time-Series Transformer (CT-PatchTST). Fig. 1 provides an overview of the model architecture. CT-PatchTST builds upon the strengths of PatchTST by introducing channel attention mechanisms that effectively capture and integrate inter-channel dependencies while retaining the benefits of channel-independent modeling.
@@ I Introduction
The key contributions of this paper are as follows:
@@ II Related Work
Time-series forecasting has long been a critical area of study, evolving from traditional statistical models to sophisticated deep learning techniques. Early approaches, such as the AutoRegressive Integrated Moving Average (ARIMA) model, laid the groundwork for this field [29]. ARIMA’s design effectively handles stationary and non-stationary data by incorporating autoregressive, moving average, and differencing components. While powerful for short-term predictions, ARIMA falls short in addressing long-term dependencies and complex non-linear relationships.
@@
Exponential smoothing methods, like the Holt-Winters model, have also been extensively used for capturing trends and seasonality. By assigning diminishing weights to older observations, these models perform well for data with seasonal variations. However, their inability to model multivariate and non-linear patterns limits their application. Gaussian Process Regression (GPR), a probabilistic framework, offers another perspective by modeling time series as distributions over functions, providing both predictions and uncertainty estimates. However, GPR struggles to scale efficiently with large datasets. Related probabilistic models like Naive Bayes and Bayesian Networks have also been used in fields such as vehicle insurance, offering high accuracy and interpretability [30].
@@
To overcome the limitations of traditional methods, deep learning models have emerged as powerful tools for time-series forecasting. Among these, Long Short-Term Memory (LSTM) networks are particularly notable for their ability to model long-term dependencies. By utilizing a cell state to retain relevant information and gate mechanisms to regulate data flow, LSTMs effectively address the vanishing gradient problem [31].
@@
Building on the advancements made by LSTMs, Seq2Seq models have brought further innovation to time-series forecasting. Initially developed for natural language processing tasks, Seq2Seq models employ an encoder-decoder structure where the encoder compresses the input sequence into a context vector, and the decoder generates predictions based on this representation. This architecture enables Seq2Seq models to handle variable-length input and output sequences, making them suitable for time-series prediction tasks with diverse temporal dynamics [32]. However, their reliance on fixed-length context vectors can result in information bottlenecks, particularly for long sequences. The introduction of attention mechanisms has helped mitigate this issue by enabling the model to focus on the most relevant parts of the input sequence dynamically. Despite their computational intensity and scaling challenges, Seq2Seq models have demonstrated notable success in capturing complex temporal dependencies, paving the way for more advanced architectures.
@@
Recently, the introduction of Transformer-based architectures has revolutionized time-series forecasting [33, 34, 35, 36, 37], with the Patch Time-Series Transformer (PatchTST) emerging as a notable example. PatchTST segments input sequences into non-overlapping patches, reducing sequence length and enabling efficient parallel processing while maintaining temporal granularity. Its self-attention mechanism dynamically assesses the importance of various time steps, making it highly effective in capturing both local and global temporal patterns. This patch-based approach not only enhances scalability but also improves model generalization across diverse datasets. Compared to Seq2Seq models, although it requires careful hyperparameter tuning and considerable computational resources during training, PatchTST has demonstrated superior performance in long-horizon forecasting tasks.
@@ III-B Reversible Instance Normalization
Reversible Instance Normalization (RevIN) is designed to mitigate the distribution shift between training and testing data in time-series applications. We apply this technique to normalize the raw data before turning it into patches. This technique operates on each time-series instance $x^{(i)}$, where $x^{(i)}$ represents the sequence of values for the $i$-th feature or channel in the dataset, with $i=1,2,\dots,M$. Each $x^{(i)}$ consists of a series of temporal values $x^{(i)}_{1},x^{(i)}_{2},\dots,x^{(i)}_{L}$].
@@ III-D Channel-Time Transformer Encoder
where $\mathbf{Z}_{\text{channel}}^{m}\in\mathbb{R}^{P\times N}$ represents the features for channel $m$, and the trainable parameters $\mathbf{W}_{h}^{Q,\text{time}},\mathbf{W}_{h}^{K,\text{time}},\mathbf{W}_{h}^{V,\text{time}}\in\mathbb{R}^{P\times d_{k}}$ are used to compute the time attention output.
@@ IV-A Dataset
We evaluate our model using two datasets that comprise measured time series data of renewable energy production, electricity consumption, and energy exchange with neighboring countries and continents, all recorded at an hourly resolution [38]. The datasets span the period from 2014 to 2019 and contain a total of 499,993 entries. Specifically, we focus on predicting three key variables: ’OffshoreWindPower,’ ’OnshoreWindPower,’ and ’SolarPowerProd.’ These variables provide valuable insights into renewable energy production patterns, making the dataset a robust benchmark for assessing the performance of our model in multivariate time series forecasting tasks.
@@
Results As shown in Table II, our model achieves the best performance compared to other methods. This result shows that applying attention to the channel data before processing the multivariate time series helps the model better learn how the different channels relate to each other. By focusing on the most important parts of the input data early on, the model can build a stronger understanding of the overall structure. This improved understanding leads to more accurate and reliable predictions.
TABLE II: Full Results on the Multivariate Forecasting Task for DK1 (Western Denmark) and DK2 (Eastern Denmark). Look-back window is 336 for all. Forecasting horizon $h\in\{96,192,336,720\}$. Best results are in bold, second-best are underlined.
Dataset Horizon (h) CT-PatchTST-512 CT-PatchTST PatchTST [26] Seq2Seq [32] iTransformer [34] PAttn [39] TimeXer [33] TSMixer [40]
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
DK1 96 0.0060 0.0515 0.0061 0.0504 0.0066 0.0540 0.0187 0.0561 0.0070 0.0549 0.0203 0.0893 0.0115 0.1004 0.0121 0.0694
192 0.0060 0.0505 0.0063 0.0502 0.0070 0.0548 0.0208 0.1190 0.0114 0.0767 0.0225 0.0938 0.0118 0.1007 0.0238 0.1098
336 0.0061 0.0509 0.0071 0.0532 0.0087 0.0595 0.1573 0.2270 0.0172 0.1260 0.0274 0.1019 0.0147 0.1091 0.0338 0.1354
720 0.0077 0.0579 0.0093 0.0644 0.0114 0.0683 0.1940 0.3510 0.0291 0.1730 0.0398 0.1198 0.0189 0.1109 0.0572 0.1656
DK2 96 0.0079 0.0598 0.0081 0.0591 0.0089 0.0634 0.0152 0.0536 0.0072 0.0629 0.0220 0.0947 0.0122 0.1011 0.0166 0.0711
192 0.0078 0.0583 0.0084 0.0590 0.0096 0.0643 0.0471 0.1623 0.0135 0.0809 0.0256 0.0979 0.0137 0.1020 0.0271 0.0901
336 0.0082 0.0579 0.0096 0.0624 0.0233 0.0687 0.0790 0.2717 0.0193 0.1570 0.0294 0.1057 0.0141 0.1097 0.0404 0.1274
720 0.0105 0.0678 0.0127 0.0750 0.0148 0.0804 0.4471 0.4609 0.0289 0.1960 0.0417 0.1239 0.0231 0.1206 0.0654 0.1655
@@ IV-D Ablation Study
In this section, we conducted two comprehensive ablation experiments to thoroughly examine the influence of varying patch lengths and prediction horizons on the performance of CT-PatchTST and to compare its effectiveness against the baseline PatchTST model with DK1 (Western Denmark) dataset. The first experiment focused on evaluating the impact of altering patch lengths, which were adjusted within a range from 4 to 40, to understand how the model’s ability to capture local and global patterns in multivariate time series data changes with different granularities. The second experiment explored the effect of varying prediction horizons, ranging from 96 to 720, to assess how well the models adapt to forecasting tasks of different temporal scales.